OpenStudy (superhelp101):

How many grams of KOH can be formed when 58.00 grams of K2O react with 44.00 grams of H2O? K2O + H2O yields KOH Be sure to balance the equation. 17.30 grams 69.10 grams 137.8 grams 273.3 grams

OpenStudy (superhelp101):

i am trying to solve this i got the balanced equation to be K2O+H2O yields 2KOH

OpenStudy (cwrw238):

K20 + H20 --> 2KOH is correct

OpenStudy (superhelp101):

yep so what do I do then?

OpenStudy (cwrw238):

work out the molecular masses of each compound from memory K = 39, O = 16 and H = 1 so 94 g of K2O react with 18 grams of water to give 112 g of KOH

OpenStudy (superhelp101):

sorry i am so confused is this right mol X g/mol= g

OpenStudy (cwrw238):

you have excess of water so 58 g of K2O will yield 58 * 112 ------ grams of KOH 94

OpenStudy (cwrw238):

theres no need to work in moles - just use grams

OpenStudy (superhelp101):

oh okay. where did you get 112 and 94?

OpenStudy (cwrw238):

best to check on those figures i gave you - in case my memory is faulty

OpenStudy (cwrw238):

K2O = 2*39 + 16 = 94 2KOH = 2 * (39 + 16 + 1) = 112

OpenStudy (superhelp101):

ooo alright so the answer is B!

OpenStudy (cwrw238):

right

OpenStudy (superhelp101):

thank you!!!

OpenStudy (cwrw238):

yw